# A Quasi-local, Functional Analytic Detection Method for Stationary Limit Surfaces of Black Hole Spacetimes

### v2

@@ Abstract
ABSTRACT.   We present a quasi-local, functional analytic method to locate and invariantly characterize the stationary limit surfaces of black hole spacetimes with stationary regions. The method is based on ellipticity-hyperbolicity transitions of the Dirac, Klein–Gordon, Maxwell, and Fierz–Pauli Hamiltonians defined on spacelike hypersurfaces of such black hole spacetimes, which occur only at the locations of stationary limit surfaces and can be ascertained from the behaviors of the principal symbols of the Hamiltonians. Therefore, since it relates solely to the effects that stationary limit surfaces have on the time evolutions of the corresponding elementary fermions and bosons, this method is profoundly different from the usual detection procedures that employ either scalar polynomial curvature invariants or Cartan invariants, which, in contrast, make use of the local geometries of the underlying black hole spacetimes. As an application, we determine the locations of the stationary limit surfaces of the Kerr–Newman, Schwarzschild–de Sitter, and Taub–NUT black hole spacetimes. Finally, we show that for black hole spacetimes with static regions, our functional analytic method serves as a quasi-local event horizon detector and gives rise to a relational concept of black hole entropy.
@@ I Introduction
This is because the usual notions of a black hole event horizon and of an ergosurface are global, observer-independent concepts that are based on particular conformal infinities of the black hole spacetime, and are thus teleological in nature (see, e.g., Booth ). To be more precise, the (outer) black hole event horizon $\mathfrak{H}$ of an asymptotically flat black hole spacetime $(\mathfrak{M},\boldsymbol{g})$ may be defined as the future boundary of the chronological past of future null infinity $\partial I^{-}(\mathscr{I}^{+})\cap\mathfrak{M}$, that is, the future boundary of the region from which one can reach future null infinity along a future-directed timelike curve Penrose2 . (For equivalent notions and for a definition of an event horizon in a black hole spacetime that is not asymptotically flat see GibbonsHawking .) The ergosurface $\mathcal{E}$ of an asymptotically flat black hole spacetime, on the other hand, is determined by the set $\{\boldsymbol{x}\in\mathfrak{M}\,|\,\boldsymbol{g}(\boldsymbol{K},\boldsymbol{K})_{|\boldsymbol{x}}=0\}$ complying with the usual condition for stationary limits, where $\boldsymbol{K}$ is a unique time-translational Killing vector field that is timelike near spacelike infinity $i^{0}$.
@@
To avoid the problem of localization that comes along with such global concepts, one can instead of the former event horizon concept employ the notion of a quasi-local black hole horizon, for instance, a trapping horizon Hayward , an isolated horizon ABDFKLW ; AshtekarKrishnan2 , or a dynamical horizon AshtekarKrishnan ; AshtekarKrishnan2 , and the associated quasi-local detection methods. These quasi-local black hole horizons preserve all essential and unproblematic features of the global black hole event horizon concepts, but they require neither the knowledge of spacetime as a whole nor the condition of asymptotic flatness. Accordingly, they can be used to quasi-locally characterize and distinguish black holes. (For a comprehensive review on the various notions of quasi-local black hole horizons, we again refer to Booth .) The typical methods for the detection of these quasi-local black hole horizons evaluate the zeros of certain scalar polynomial curvature invariants AbdelqaderLake ; Lake ; McNuttPage ; PageShoom ; TavlayanTekin or of Cartan invariants BrooksEtAl , as these invariants are nonzero away from the horizons but vanish on the horizons. (Traversing from the outside to the inside of these surfaces, they change their sign.) The scalar polynomial curvature invariants found in KarlhedeEtAl ; GassEtAl , on the other hand, vanish on stationary limit surfaces, which enclose regions of spacetime where no observer can remain at rest. Thus, aside from the event horizons of black hole spacetimes with static regions, they can also be used to detect the ergosurfaces of black hole spacetimes with stationary, rotational regions. The field of main application of the above invariants is to be found in numerical general relativity, where they are frequently employed as tools to track quasi-local horizons in the numerical evolution of black holes that have settled down to becoming either stationary or near-stationary, providing the horizons’ exact or approximate locations on spacelike slices of the black hole spacetimes, respectively (see, e.g., PetersColeySchnetter for an application to black hole mergers).
@@
In the present paper, we introduce a quasi-local, functional analytic method to detect the stationary limit surfaces of black hole spacetimes with stationary regions. [^1] A region $\mathcal{U}\subseteq\mathfrak{M}$ of a spacetime $(\mathfrak{M},\boldsymbol{g})$ is called stationary if it admits a unique time-translational Killing vector field $\boldsymbol{K}$ that is timelike everywhere in $\mathcal{U}$, i.e., $\boldsymbol{g}(\boldsymbol{K},\boldsymbol{K})_{|\boldsymbol{x}}>0$ for all $\boldsymbol{x}\in\mathcal{U}$. This detection method is different from the above quasi-local, differential geometric detection methods in so far as, instead of the local geometries of the black hole spacetimes, it relies on the effects that stationary limit surfaces have on the time evolutions of the elementary spin-$\tfrac{1}{2}$ fermions and the elementary spin-00, spin-$1$, and spin-$2$ bosons. The method is based on the observation that the Dirac, Klein–Gordon, Maxwell, and Fierz–Pauli Hamiltonians on spacelike hypersurfaces of black hole spacetimes with stationary regions undergo ellipticity-hyperbolicity transitions precisely at the locations of stationary limit surfaces, such as the outer and inner ergosurfaces of the Kerr–Newman black hole spacetime. (For the definitions of the different Hamiltonians and the notions of ellipticity and hyperbolicity see Section II.) Accordingly, one can locate stationary limit surfaces of black hole spacetimes with stationary regions by analyzing the determinantial polynomials of the principal symbols of these Hamiltonians. This is worked out in detail in Section III, where, as an application, we also use our detection method to locate the stationary limit surfaces of the Kerr–Newman, Schwarzschild–de Sitter, and Taub–NUT black hole spacetimes. We point out that the functional analytic detection method has the advantage that its computational complexity in numerical algorithms is in general substantially lower compared to those of the differential geometric detection methods as, instead of an (at most) second-order system of partial differential equations, only a quadratic polynomial has to be evaluated. Finally, since for black hole spacetimes with static regions the stationary limit surfaces coincide with the event horizons, our method can be employed, at least in this case, to quasi-locally detect event horizons. This, as well as a relational concept of black hole entropy for asymptotically flat black hole spacetimes with static regions, is discussed in Section IV.
@@ II Preliminaries
We first specify the differential geometric setting. We let $(\mathfrak{M},\boldsymbol{g})$ be a smooth, connected, oriented and time-oriented Lorentzian $4$-manifold with metric signature $(1,3,0)$ and $\mathcal{V}$ and $\mathcal{W}$ finite rank $\mathbb{K}$-vector bundles with base space $\mathfrak{M}$ and fiberwise metrics $\boldsymbol{g}_{\mathcal{V}}$ and $\boldsymbol{g}_{\mathcal{W}}$, where $\mathbb{K}=\mathbb{R}\,\,\textnormal{or}\,\,\mathbb{C}$. Moreover, we let $\boldsymbol{\nabla}$ be the Levi–Civita connection on $T\mathfrak{M}$ and $\boldsymbol{\nabla}_{\mathcal{V}}$ and $\boldsymbol{\nabla}_{\mathcal{W}}$ the connections on $\mathcal{V}$ and $\mathcal{W}$, respectively. We assume that $(\mathfrak{M},\boldsymbol{g})$ is stably causal, i.e., it admits a foliation by a family of spacelike hypersurfaces $(\mathfrak{N}_{t})_{t\in\mathbb{R}}$, with $\mathfrak{N}_{t}:=\{t\}\times\mathfrak{N}$, where the label $t$ is a global temporal function on $\mathfrak{M}$ MinguzziSanchez . This assumption implies that $\mathfrak{M}$ has the product structure $\mathfrak{M}=\mathbb{R}\times\mathfrak{N}$. Consequently, if $\mathfrak{N}$ is complete, $(\mathfrak{M},\boldsymbol{g})$ is globally hyperbolic.
@@
A linear, matrix-valued differential operator of order (at most) $m$ on $\mathfrak{M}$ is a mapping $P:\Gamma^{\infty}(\mathfrak{M},\mathcal{V})\rightarrow\Gamma^{\infty}(\mathfrak{M},\mathcal{W})$ that in any local coordinate system $\boldsymbol{x}$ on an open subset $\mathcal{U}\subseteq\mathfrak{M}$ and any local trivializations $\mathcal{V}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{K}^{p}$ and $\mathcal{W}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{K}^{q}$, with $(p,q)\in\mathbb{N}^{2}$, is of the form
@@
The principal symbol of a linear, matrix-valued differential operator $P\in\textnormal{Diff}_{m}(\mathcal{V},\mathcal{W})$ is a bundle homomorphism $\sigma_{P}:\mathcal{V}_{\boldsymbol{x}}\rightarrow\mathcal{W}_{\boldsymbol{x}}$ that in any local coordinate system $\boldsymbol{x}$ on an open subset $\mathcal{U}\subseteq\mathfrak{M}$ and for any $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{x}}\mathfrak{M}$ reads
@@
and a Clifford multiplication $\boldsymbol{\gamma}:T\mathfrak{M}\rightarrow\textnormal{End}(S\mathfrak{M})$ that is fiber-preserving and satisfies the anticommutation relations
$\{\boldsymbol{\gamma}(\boldsymbol{v}),\boldsymbol{\gamma}(\boldsymbol{w})\}=2\hskip 1.42271pt\boldsymbol{g}(\boldsymbol{v},\boldsymbol{w})\,\mbox{\rm 1 1}_{S_{\boldsymbol{x}}\mathfrak{M}}$ (5)
for all $\boldsymbol{x}\in\mathfrak{M}$ and $\boldsymbol{v},\boldsymbol{w}\in T_{\boldsymbol{x}}\mathfrak{M}$. The matrix $\mathscr{S}$ in the indefinite inner product (4) is chosen such that $\boldsymbol{\gamma}^{\dagger}\circ\mathscr{S}=\mathscr{S}\circ\boldsymbol{\gamma}$. Then, the Dirac operator $\mathcal{D}:\Gamma^{\infty}(\mathfrak{M},S\mathfrak{M})\rightarrow\Gamma^{\infty}(\mathfrak{M},S\mathfrak{M})$ with an external potential $\mathcal{B}\in C^{\infty}\bigl{(}\mathfrak{M},\textnormal{Mat}(4\times 4,\mathbb{C})\bigr{)}$ in any local coordinate system $\boldsymbol{x}$ on an open subset $\mathcal{U}\subseteq\mathfrak{M}$ and in any local trivialization $S\mathfrak{M}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{C}^{4}$ takes the form
where $\nabla_{\mu}$ are the components of the covariant derivative corresponding to the unique metric connection on $S\mathfrak{M}$ obtained as a lift of the Levi–Civita connection $\boldsymbol{\nabla}$ on $T\mathfrak{M}$. Using this operator, we can define the Dirac equation as
@@
in which the Dirac wave function $\boldsymbol{\psi}\in\Gamma^{\infty}_{\textnormal{sc}}(\mathfrak{M},S\mathfrak{M})$, the subscript “sc” denoting spatially compact support, and $m\in\mathbb{R}_{\geq 0}$ is the rest mass of the spin-$\tfrac{1}{2}$ fermion associated with the Dirac wave function. Choosing local coordinates $\boldsymbol{x}=(t,\boldsymbol{y})$ on ×$\mathbb{R}\times\mathcal{S}\subseteq\mathfrak{M}$ and multiplying this equation by the inverse of $\gamma^{t}$, separation of the partial time derivative term yields the Dirac equation in Hamiltonian form
@@
where $\mathcal{H}_{\mathcal{D}}:\Gamma^{\infty}(\mathfrak{N},S\mathfrak{M})\rightarrow\Gamma^{\infty}(\mathfrak{N},S\mathfrak{M})$ is the Dirac Hamiltonian and $\omega_{t\mu\nu}$ are the temporal components of the connection $1$-forms $\omega_{\mu\nu}=\boldsymbol{g}(\boldsymbol{\nabla}\boldsymbol{e}_{\mu},\boldsymbol{e}_{\nu})$ of the Levi–Civita connection with respect to a local orthonormal frame $(\boldsymbol{e}_{\mu})_{\mu=0,...,3}$. For time-independent spacetimes, the Dirac Hamiltonian with domain of definition $\textnormal{dom}(\mathcal{H}_{\mathcal{D}})=\Gamma^{\infty}_{\textnormal{c}}(\mathfrak{N}\backslash\partial\mathfrak{N},S\mathfrak{M})$, the subscript “c” signifying compact support, is formally self-adjoint with respect to the scalar product
@@
In order to define the Klein–Gordon, Maxwell, and Fierz–Pauli Hamiltonians, which determine the time evolutions of the elementary spin-00, spin-$1$, and spin-$2$ bosons, respectively, we employ the d’Alembert operator $\square:\Gamma^{\infty}(\mathfrak{M},\mathfrak{M}\times\mathbb{K})\rightarrow\Gamma^{\infty}(\mathfrak{M},\mathfrak{M}\times\mathbb{K})$ with an external potential $\mathfrak{B}\in C^{\infty}(\mathfrak{M},\mathbb{K})$ that in any local coordinate system $\boldsymbol{x}$ on $\mathcal{U}$ and in any local trivialization $\mathfrak{M}\times\mathbb{K}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{K}$ reads
@@
where the scalar wave function $\phi\in\Gamma^{\infty}_{\textnormal{sc}}(\mathfrak{M},\mathfrak{M}\times\mathbb{C})$ and $\mathfrak{m}\in\mathbb{R}_{\geq 0}$ is the rest mass of the spin-00 boson associated with the scalar wave function, the (homogeneous) electromagnetic wave equation as
@@
$i\partial_{t}\Phi(t,\boldsymbol{y})=\frac{1}{g^{tt}}\begin{pmatrix}0&g^{tt}\\ \varepsilon&-i\kappa\end{pmatrix}\Phi(t,\boldsymbol{y})=:\mathcal{H}_{\textnormal{KG}}\Phi(t,\boldsymbol{y})\,,$ (10)
with the Klein–Gordon Hamiltonian $\mathcal{H}_{\textnormal{KG}}:\Gamma^{\infty}(\mathfrak{N},\mathfrak{M}\times\mathbb{C})\rightarrow\Gamma^{\infty}(\mathfrak{N},\mathfrak{M}\times\mathbb{C})$, the wave function $\Phi(t,\boldsymbol{y}):=(1,i\partial_{t})^{\textnormal{T}}\hskip 0.56917pt\phi(t,\boldsymbol{y})$, and
@@
[…]

### v3

@@ Abstract
ABSTRACT.   We present a quasi-local, functional analytic method to locate and invariantly characterize the stationary limit surfaces of black hole spacetimes with stationary regions. The method is based on ellipticity-hyperbolicity transitions of the Dirac, Klein–Gordon, Maxwell, and Fierz–Pauli Hamiltonians defined on spacelike hypersurfaces of such black hole spacetimes, which occur only at the locations of stationary limit surfaces and can be ascertained from the behaviors of the principal symbols of the Hamiltonians. Therefore, since it relates solely to the effects that stationary limit surfaces have on the time evolutions of the corresponding elementary fermions and bosons, this method is profoundly different from the usual detection procedures that employ either scalar polynomial curvature invariants or Cartan invariants, which, in contrast, make use of the local geometries of the underlying black hole spacetimes. As an application, we determine the locations of the stationary limit surfaces of the Kerr–Newman, Schwarzschild–de Sitter, and Taub–NUT black hole spacetimes. Finally, we show that for black hole spacetimes with static regions, our functional analytic method serves as a quasi-local event horizon detector and gives rise to relational concepts of event horizons and black hole entropy.
@@ I Introduction
This is because the usual notions of a black hole event horizon and of an ergosurface are global, observer-independent concepts that are based on conformal infinity of the black hole spacetime, and are thus teleological in nature (see, e.g., Booth ). To be more precise, the (outer) black hole event horizon $\mathfrak{H}$ of an asymptotically flat black hole spacetime $(\mathfrak{M},\boldsymbol{g})$ may be defined as the future boundary of the chronological past of future null infinity $\partial I^{-}(\mathscr{I}^{+})\cap\mathfrak{M}$, that is, the future boundary of the region from which one can reach future null infinity along a future-directed, timelike curve Penrose2 . (For equivalent notions and for a definition of an event horizon in a black hole spacetime that is not asymptotically flat see GibbonsHawking .) The ergosurface $\mathcal{E}$ of an asymptotically flat black hole spacetime, on the other hand, is determined by the set $\bigl\{\boldsymbol{x}\in\mathfrak{M}\,\big|\,\boldsymbol{g}(\boldsymbol{K},\boldsymbol{K})_{|\boldsymbol{x}}=0\bigr\}$ complying with the usual condition for stationary limits, where $\boldsymbol{K}$ is a unique time-translational Killing vector field that is timelike near spacelike infinity $i^{0}$.
@@
To avoid the problem of localization that comes along with such global concepts, one can instead of, e.g., the former event horizon concept employ the notion of a quasi-local black hole horizon, for instance, a trapping horizon Hayward , an isolated horizon ABDFKLW ; AshtekarKrishnan2 , or a dynamical horizon AshtekarKrishnan ; AshtekarKrishnan2 , and the associated quasi-local detection methods. These quasi-local black hole horizons preserve all essential and unproblematic features of the global black hole event horizon concepts, but they require neither the knowledge of spacetime as a whole nor the condition of asymptotic flatness. Accordingly, they can be used to quasi-locally characterize and distinguish black holes. (For a comprehensive review on the various notions of quasi-local black hole horizons, we again refer to Booth .) The typical methods for the detection of these quasi-local black hole horizons evaluate the zeros of certain scalar polynomial curvature invariants AbdelqaderLake ; Lake ; McNuttPage ; PageShoom ; TavlayanTekin or of Cartan invariants BrooksEtAl , as these invariants are nonzero away from the horizons but vanish on the horizons. (Traversing from the outside to the inside of these surfaces, they change their sign.) The scalar polynomial curvature invariants found in KarlhedeEtAl ; GassEtAl , on the other hand, vanish on stationary limit surfaces. Thus, aside from the event horizons of black hole spacetimes with static regions, they can also be used to locally detect the ergosurfaces of black hole spacetimes with stationary, rotational regions. The main field of application of the above invariants is to be found in numerical relativity, where they are frequently employed as tools to track quasi-local horizons in the numerical evolution of black holes that have settled down to becoming either stationary or near-stationary, providing the horizons’ exact or approximate locations on spacelike slices of the black hole spacetimes, respectively (see, e.g., PetersColeySchnetter for an application to black hole mergers).
@@
In the present paper, we introduce a quasi-local, functional analytic method to detect the stationary limit surfaces of black hole spacetimes with stationary regions. [^1] A region $\mathcal{U}\subseteq\mathfrak{M}$ of a spacetime $(\mathfrak{M},\boldsymbol{g})$ is called stationary if it admits a unique time-translational Killing vector field $\boldsymbol{K}$ that is timelike everywhere in $\mathcal{U}$, i.e., $\boldsymbol{g}(\boldsymbol{K},\boldsymbol{K})_{|\boldsymbol{x}}>0$ for all $\boldsymbol{x}\in\mathcal{U}$. This detection method is different from the above quasi-local, differential geometric detection methods in so far as, instead of the local geometries of the black hole spacetimes, it relies on the effects that stationary limit surfaces have on the time evolutions of elementary spin-$\tfrac{1}{2}$ fermions and elementary spin-0, spin-11, and spin-$2$ bosons. The method is based on the observation that the Dirac, Klein–Gordon, Maxwell, and Fierz–Pauli Hamiltonians on spacelike hypersurfaces of black hole spacetimes with stationary regions undergo ellipticity-hyperbolicity transitions precisely at the locations of stationary limit surfaces, such as the outer and inner ergosurfaces of the Kerr–Newman black hole spacetime. (For the definitions of the different Hamiltonians and the notions of ellipticity and hyperbolicity see Section II.) Accordingly, one can locate stationary limit surfaces of black hole spacetimes with stationary regions by analyzing the determinantial polynomials of the principal symbols of these Hamiltonians. This is worked out in detail in Section III, where, as an application, we also use our detection method to locate the stationary limit surfaces of the Kerr–Newman, Schwarzschild–de Sitter, and Taub–NUT black hole spacetimes. We point out that the functional analytic detection method has the advantage that its computational complexity in numerical algorithms is in general substantially lower compared to those of the differential geometric detection methods as, instead of an (at most) fourth-order nonlinear system of partial differential equations, only a quadratic polynomial of three indeterminates with possibly transcendental algebraic conditions on its coefficients has to be evaluated. Finally, since for black hole spacetimes with static regions the stationary limit surfaces coincide with the event horizons, our method can be employed, at least in this case, to quasi-locally detect event horizons. This, as well as relational concepts of event horizons and black hole entropy for black hole spacetimes with static regions, is discussed in Section IV.
@@ II Preliminaries
[…]
@@
A linear, matrix-valued differential operator of order (at most) $m$ on $\mathfrak{M}$ is a mapping $P\hskip-1.9919pt:\Gamma^{\infty}(\mathfrak{M},\mathcal{V})\rightarrow\Gamma^{\infty}(\mathfrak{M},\mathcal{W})$ that in any local coordinate system $\boldsymbol{x}$ on an open subset $\mathcal{U}\subseteq\mathfrak{M}$ and any local trivializations $\mathcal{V}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{K}^{p}$ and $\mathcal{W}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{K}^{q}$, with $(p,q)\in\mathbb{N}^{2}$, is of the form
@@
The principal symbol of a linear, matrix-valued differential operator $P\in\textnormal{Diff}_{m}(\mathcal{V},\mathcal{W})$ is a bundle homomorphism $\sigma_{P}\hskip-1.9919pt:\mathcal{V}_{\boldsymbol{x}}\rightarrow\mathcal{W}_{\boldsymbol{x}}$ that in any local coordinate system $\boldsymbol{x}$ on an open subset $\mathcal{U}\subseteq\mathfrak{M}$ and for any $\boldsymbol{\xi}\in T^{\star}_{\boldsymbol{x}}\mathfrak{M}$ reads
@@
$\langle\,.\,|\,.\,\rangle_{\boldsymbol{x}}\hskip-1.9919pt:S_{\boldsymbol{x}}\mathfrak{M}\times S_{\boldsymbol{x}}\mathfrak{M}\rightarrow\mathbb{C}\,,\quad(\boldsymbol{\psi},\boldsymbol{\chi})\mapsto\boldsymbol{\psi}^{\dagger}\mathscr{S}\hskip 0.56917pt\boldsymbol{\chi}\,$, (4)
and a Clifford multiplication $\boldsymbol{\gamma}\hskip-1.9919pt:T_{\boldsymbol{x}}\mathfrak{M}\rightarrow\textnormal{End}(S_{\boldsymbol{x}}\mathfrak{M})$ that is fiber-preserving and satisfies the anticommutation relations
$\bigl\{\boldsymbol{\gamma}(\boldsymbol{v}),\boldsymbol{\gamma}(\boldsymbol{w})\bigr\}=2\hskip 1.42271pt\boldsymbol{g}(\boldsymbol{v},\boldsymbol{w})\,\mbox{\rm 1 \hskip-10.50005pt 1}_{S_{\boldsymbol{x}}\mathfrak{M}}$ (5)
for all $\boldsymbol{x}\in\mathfrak{M}$ and $\boldsymbol{v},\boldsymbol{w}\in T_{\boldsymbol{x}}\mathfrak{M}$. The matrix $\mathscr{S}$ in the indefinite inner product (4) is chosen such that $\boldsymbol{\gamma}^{\dagger}\circ\mathscr{S}=\mathscr{S}\circ\boldsymbol{\gamma}$. Then, the Dirac operator $\mathcal{D}\hskip-1.9919pt:\Gamma^{\infty}(\mathfrak{M},S\mathfrak{M})\rightarrow\Gamma^{\infty}(\mathfrak{M},S\mathfrak{M})$ with an external potential $\mathcal{B}\in\textnormal{End}(S\mathfrak{M})$ in any local coordinate system $\boldsymbol{x}$ on an open subset $\mathcal{U}\subseteq\mathfrak{M}$ and in any local trivialization $S\mathfrak{M}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{C}^{4}$ takes the form
@@
in which the Dirac wave function $\boldsymbol{\psi}\in\Gamma^{\infty}_{\textnormal{sc}}(\mathfrak{M},S\mathfrak{M})$, the subscript “sc” denoting spatially compact support, and $m\in\mathbb{R}_{\geq 0}$ is the rest mass of the spin-$\tfrac{1}{2}$ fermion associated with the Dirac wave function. Choosing local coordinates $\boldsymbol{x}=(t,\boldsymbol{y})$ on $\mathbb{R}\times\mathcal{S}\subseteq\mathfrak{M}$ and multiplying this equation by the inverse of $\gamma^{t}$, separation of the partial time derivative term yields the Dirac equation in Hamiltonian form
@@
where $\mathcal{H}_{\mathcal{D}}\hskip-1.9919pt:\Gamma^{\infty}(\mathfrak{N},S\mathfrak{M})\rightarrow\Gamma^{\infty}(\mathfrak{N},S\mathfrak{M})$ is the Dirac Hamiltonian and $\omega_{t\mu\nu}$ are the temporal components of the connection $1$-forms $\omega_{\mu\nu}=\boldsymbol{g}(\boldsymbol{\nabla}\boldsymbol{e}_{\mu},\boldsymbol{e}_{\nu})$ of the Levi–Civita connection with respect to a local orthonormal frame $(\boldsymbol{e}_{\mu})_{\mu\in\{0,1,2,3\}}$. For time-independent spacetimes, the Dirac Hamiltonian with domain of definition $\textnormal{Dom}(\mathcal{H}_{\mathcal{D}})=\Gamma^{\infty}_{\textnormal{c}}(\mathfrak{N}\backslash\partial\mathfrak{N},S\mathfrak{M})$, the subscript “c” signifying compact support, is formally self-adjoint with respect to the scalar product
@@
In order to define the Klein–Gordon, Maxwell, and Fierz–Pauli Hamiltonians, which determine the time evolutions of the elementary spin-0, spin-11, and spin-$2$ bosons, respectively, we employ the d’Alembert operator $\square\hskip-1.9919pt:\Gamma^{\infty}(\mathfrak{M},\mathfrak{M}\times\mathbb{K})\rightarrow\Gamma^{\infty}(\mathfrak{M},\mathfrak{M}\times\mathbb{K})$ with an external potential $\mathfrak{B}\in\textnormal{End}(\mathfrak{M}\times\mathbb{K})$ that in any local coordinate system $\boldsymbol{x}$ on $\mathcal{U}$ and in any local trivialization $\mathfrak{M}\times\mathbb{K}_{|\mathcal{U}}\rightarrow\mathcal{U}\times\mathbb{K}$ reads
@@
where the scalar wave function $\phi\in\Gamma^{\infty}_{\textnormal{sc}}(\mathfrak{M},\mathfrak{M}\times\mathbb{C})$ and $\mathfrak{m}\in\mathbb{R}_{\geq 0}$ is the rest mass of the spin-0 boson associated with the scalar wave function, the (homogeneous) electromagnetic wave equation as
@@
$i\partial_{t}\Phi(t,\boldsymbol{y})=\frac{1}{g^{tt}}\hskip-1.42271pt\begin{pmatrix}0&g^{tt}\\ \varepsilon&-i\kappa\end{pmatrix}\hskip-1.9919pt\Phi(t,\boldsymbol{y})=:\mathcal{H}_{\textnormal{KG}}\Phi(t,\boldsymbol{y})\,$, (10)
with the Klein–Gordon Hamiltonian $\mathcal{H}_{\textnormal{KG}}\hskip-1.9919pt:\Gamma^{\infty}(\mathfrak{N},\mathfrak{M}\times\mathbb{C})\rightarrow\Gamma^{\infty}(\mathfrak{N},\mathfrak{M}\times\mathbb{C})$, the wave function $\Phi(t,\boldsymbol{y}):=(1,i\partial_{t})^{\textnormal{T}}\hskip 0.56917pt\phi(t,\boldsymbol{y})$, and
@@
 $\mbox{$\prec\!\!$ \nolinebreak}\Phi\hskip 0.56917pt|\hskip 0.56917pt\Psi\mbox{\nolinebreak$\succ$}_{\mathfrak{N}}=\int_{\mathfrak{N}}\Phi^{\dagger}\hskip-1.9919pt\begin{pmatrix}\varepsilon/g^{tt}&0\\ 0&1\end{pmatrix}\hskip-1.42271pt\Psi\,\textnormal{d}\mu_{\mathfrak{N}}(\boldsymbol{x})\,.$